4-43 Overview of general ledger relationships. Estevez Company uses normal costing in its job-costing system. The company produces kitchen cabinets. The beginning balances (December 1) and ending bal- ances (as of December 30) in their inventory accounts are as follows: Beginning Balance 12/1 Ending Balance 12/30 Materials Control Work-in-Process Control Manufacturing Department Overhead Control Finished Goods Control S 17,000 8,000 88,000 38,800 $ 4,200 3,400 8,800
Answer #1
1. JEs for Dec.31 payroll Debit Credit
WIP control 8600
Mfg. dept. OH control 2800
Cash 11400
2.. Total amount of materials requisitioned into WIP during Dec.
Materials control a/c
Beginning balance 4200
Purchases during Dec 132600
Materials requisitioned into WIP during Dec.(plug-in) --------------------(ANSWER) 119800
Ending Balance 17000
136800 136800
2.b.Amt. of direct mfg. labor recorded in WIP during Dec
Cost of goods manufactured for Dec(Given) 468000
Materials requisitioned (as found in 2.a.) 119800
Mfg.labor                                 (suppose) (X+8600)
Mfg.OH applied        (Given 110% *D/L cost) 110%*(x+8600)
So,forming an equation,
468000=119800+(x+8600)+(1.1*(x+8600))
Total Mfg. Labor, for Dec(157210+8600)= 165810
2. c.Total Amt. of mfg. OH recorded in WIP during Dec
165810*110%
182391
d..WIP control a/c
Beg. Balance 13400
Materials requisitioned into WIP during Dec. 119800
MFG.   Labor               (as in 2.b.) 165810
Applied mfg. OH        (as in 2.c.) 182391
Cost of goods manufactured 468000
Ending balance         (ANSWER) 13401
481401 481401
e.Cost of Goods Sold for Dec.
Op.Finished goods 8800
Add: COG mfd. 468000
Less: Cl.fin.gds. -38800
COGS 438000
3..
3...Actual MFG. OH incurred for Dec.(188000+2800)= 190800
Mfg. OH applied as in WIP control a/c 182391
MFG. OH under-applied 8409
Mfg. OH control a/c
Incurred till dec 30 188000
Dec 31,Ind.labor 2800
Applied to WIP 182391
Under-applied OH trfd. To COGS 8409
190800 190800
Journal Entry
COGS 8409
Mfg. OH control a/c 8409
COGS= 438000+8409= 446409
Or 446410